# Closed-Form Feedback-Free Learning with Forward Projection

### v3

@@ A.9 Optimisation Theory
$\mathbb{E}\left[\frac{\lVert\textbf{Y}-\textbf{P}_{\textbf{a}}\textbf{Y}\rVert_{F}^{2}}{\lVert\textbf{Y}\rVert_{F}^{2}}\right]<\frac{N-m}{N}$ (45)
Thus, even under pure Gaussian noise, the two-layer model achieves a lower expected error than the simple linear predictor.
@@ A.10 Iterative learning with Forward Projection
where $\mathbf{z}_{l}^{(j)}=\mathbf{a}_{l-1}^{(j)}\mathbf{W}_{l}$. The gradient of $\mathcal{L}_{l}^{(t)}$ with respect to $\mathbf{W}_{l}$ is:
@@
where $\eta$ is the learning rate. Optionally, regularisation may be included via the ridge penalty:
$\nabla_{\mathbf{W}_{l}}\mathcal{L}_{l}^{(t)}\;+=\;2\lambda\mathbf{W}_{l}.$ (50)
@@
A.11 Modelling high dimensional labels
From (28), we have:
$\hat{\textbf{y}}_{l}:=g_{l}^{-1}\left(\textbf{z}_{l}-g_{l}(\textbf{a}_{l-1}\textbf{Q}_{l})\right)\textbf{U}_{l}^{+}.$ (51)
@@ A.11 Modelling high dimensional labels
$\hat{\textbf{y}}_{l}:=\left(\textbf{z}_{l}-\textbf{a}_{l-1}\textbf{Q}_{l}\right)\textbf{U}_{l}^{+}.$ (52)
@@
Recalling that the realised potentials are estimates of pre-defined target potentials, such that $\textbf{z}_{l}=\tilde{\textbf{z}}_{l}-\textbf{e}_{l}$, we have
$\hat{\textbf{y}}_{l}:=\left(\tilde{\textbf{z}}_{l}-\textbf{e}_{l}-\textbf{a}_{l-1}\textbf{Q}_{l}\right)\textbf{U}_{l}^{+}.$ (53)
Applying the identity function as $g_{l}$ in equation (4), we have
$\tilde{\textbf{z}}_{l}=\textbf{a}_{l-1}\textbf{Q}_{l}+\textbf{y}\textbf{U}_{l}.$ (54)
Substituting, we have
$\hat{\textbf{y}}_{l}=\left(\textbf{y}\textbf{U}_{l}-\textbf{e}_{l}\right)\textbf{U}_{l}^{+}.$ (55)
Assuming, that U has full row rank, and that $m_{l}>m_{L}$, we have $\textbf{U}_{l}\textbf{U}^{+}_{l}=\textbf{I}_{m_{L}}$. Therefore
$\hat{\textbf{y}}_{l}=\textbf{y}-\textbf{e}_{l}\textbf{U}_{l}^{+}.$ (56)
Assuming $\textbf{e}_{l}\sim\mathcal{N}(0,\epsilon\textbf{I}_{m_{l}})$ for some $\epsilon\in\mathbb{R}$, we have $\mathbb{E}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]=0$, as U is random and independent of $\textbf{e}_{l}$. Therefore
$\mathbb{E}\left[\hat{\textbf{y}}_{l}\right]=\textbf{y}-\mathbb{E}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]=\textbf{y}.$ (57)
Therefore $\operatorname{Var}\left[\tilde{\textbf{y}}_{l}-\textbf{y}_{l}\right]=\operatorname{Var}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]$. Using the property of linear transformation of Gaussian variables, we have
$\operatorname{Var}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]=\epsilon\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}.$ (58)
We note that
$\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)=\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}$ (59)
Furthermore, it is noted that
$\mathbb{E}\left[\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}\right]=\mathbb{E}\left[\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)\right]\approx\mathbb{E}[\operatorname{Tr}\left((\textbf{U}_{l}\textbf{U}_{l}^{T})^{-1}\right)]$ (60)
We observe that the Frobenius norm $\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}$ equals the sum of inverse squared singular values
$\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}=\sum_{i=1}^{m_{L}}\frac{1}{\sigma_{i}}^{2}.$ (61)
Selecting $\textbf{U}_{l}$ with independent and identically distributed Gaussian entries, the singular values squared are the eigenvalues of the Wishart matrix
$\textbf{U}_{l}\textbf{U}_{l}^{T}\sim\mathcal{W}_{m_{L}}(m_{l},\textbf{I}_{m_{l}}).$ (62)
Therefore we have $\mathbb{E}[\operatorname{Tr}\left((\textbf{U}_{l}\textbf{U}_{l}^{T})^{-1}\right)]=\frac{m_{L}}{m_{l}-m_{L}-1}$ for $m_{L}+1<m_{l}$. Thus, the expectation $\mathbb{E}\left[\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)\right]$ increases with $m_{L}$ and decreases with $m_{l}$. Thus, higher dimensional labels are subject to poorer encoding, for a given hidden layer dimension. Conversely, for a given label dimension, it is expected that representation and learning will improve with increases of the hidden layer dimension. Accordingly, adequate modelling of high dimensional labels may require larger hidden layer dimensions.
@@ A.12 Training complexity of other methods
In this section, we describe the training complexity involved in each step of the different methods that are used in this paper for benhcmarking purposes.
@@ A.12.2 Local Supervision
In local supervision, after the forward pass through the main layer ($N_{e}Nm^{2}$ MACs), an auxiliary forward pass is conducted through the local supervision head ($N_{e}Nmm_{L}$ MACs). Subsequently, the auxiliary gradient is computed at the local supervision head ($N_{e}Nmm_{L}$ MACs), which is backpropagated through the main layer ($N_{e}Nm^{2}$ MACs). Forward weight updates for the main layer occur at the end of each batch ($N_{e}B^{-1}Nm^{2}$ MACs). Optionally, the auxiliary layer may be trained ($N_{e}B^{-1}Nmm_{L}$ MACs). Parameters are stored for the main layer weights ($m^{2}$) and auxiliary weights $mm_{L}$. Local supervision requires storage of main layer weights ($m^{2}$), auxiliary layer weights ($mm_{L}$), main layer input activations ($Bm$), main layer output activations ($Bm$), auxiliary layer output activations ($Bm_{L}$) and weight updates ($m^{2}$).
@@ A.12.3 Forward-Forward (FF)
For a given layer, FF requires two forward passes – for “positive” and “negative” data, respectively ($2N_{e}Nm^{2}$ MACs) [15]. A goodness metric is computed between positive and negative activations, requiring $N_{e}Nm$ MACs. Local gradients are computed for both positive and negative passes ($2N_{e}Nm^{2}$ MACs). Forward weight updates occur with each batch ($B^{-1}N_{e}Nm^{2}$ MACs). At inference time, forward passes are required for each possible class label ($Nm^{2}m_{l}$ MACs). Forward-Forward requires storage of weight parameters ($m^{2}$), positive and negative input activations ($2B(m+m_{L})$), positive and negative output activations ($2Bm$) and weight updates ($m^{2}$).
@@ A.12.4 Difference Target Propagation (DTP)
DTP replaces gradient transport with targets formed using a learned approximate inverse (decoder) and a difference correction. Training requires an encoder forward pass ($N_{e}Nm^{2}$ MACs), and two decoder passes ($2N_{e}Nm^{2}$ MACs) to perform the difference correction operation between targets and observed activations [1]. Decoder training employs a denoising style loss, forwarding through both encoder ($N_{e}Nm^{2}$ MACs) and decoder ($N_{e}Nm^{2}$ MACs), before computing the decoder weight gradient ($N_{e}Nm^{2}$ MACs). Finally, the encoder weight gradient is computed with respect to the local target loss ($N_{e}Nm^{2}$ MACs). Weight updates for both encoder and decoder occur after each batch ($2B^{-1}N_{e}Nm^{2}$ MACs). At inference time, only the encoder forward pass is required. Assuming batch size $B=1$, training requires a forward pass ($N_{e}Nm^{2}$ MACs) and $8N_{e}Nm^{2}$ further MACs per layer. DTP requires storage of encoder and decoder weight parameters ($2m^{2}$), input activations ($Bm$), output activations ($Bm$), targets ($Bm$) and weight updates ($m^{2}$), which can be performed sequentially for encoder and decoder.
@@ A.12.5 Predictive Coding
PC performs $k$ iterations of activity inference that alternate forward prediction ($kN_{e}Nm^{2}$ MACs) and feedback error projection through a decoder network ($kN_{e}Nm^{2}$ MACs), followed by a local Hebbian weight gradient computation ($N_{e}Nm^{2}$ MACs) [5]. Forward weight updates occur at the end of each batch ($N_{e}B^{-1}Nm^{2}$ MACs). Letting $k=1$, $B=1$, training requires a forward pass ($N_{e}Nm^{2}$ MACs) and $3N_{e}N_{m}^{2}$ further MACs per layer. Predictive coding requires storage of weight parameters ($m^{2}$), input activations ($m$), output activations ($Bm$), output targets ($Bm$) and hebbian weight updates ($m^{2}$).

### v4

@@ A.9 Optimisation Theory
$E_{\mathrm{FP}}=\mathbb{E}\left[\lVert(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\mathbf{Y}_{\perp}\rVert^{2}_{F}\right].$ (48)
As $\mathbf{P}_{\mathbf{A}}\mathbf{Y}_{\perp}$ is orthogonal to $(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\mathbf{Y}_{\perp}$, we may apply the Pythagorean theorem [49], such that
$\lVert\mathbf{Y}_{\perp}\rVert^{2}_{F}=\lVert\mathbf{P}_{\mathbf{A}}\mathbf{Y}_{\perp}\rVert^{2}_{F}+\lVert(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\mathbf{Y}_{\perp}\rVert^{2}_{F}.$ (49)
Substituting (36) and (45), we have
$E_{\mathrm{linear}}=\mathbb{E}\left[\lVert\mathbf{P}_{\mathbf{A}}\mathbf{Y}_{\perp}\rVert^{2}_{F}\right]+E_{\mathrm{FP}}.$ (50)
Therefore, under the inclusion assumption that $\operatorname{col}(\hat{\mathbf{Y}}_{\mathrm{linear}})\subset\operatorname{col}(\mathbf{A})$, the improvement of the Forward Projection model over the linear model is given by
$\boxed{E_{\mathrm{linear}}-E_{\mathrm{FP}}=\mathbb{E}\left[\lVert\mathbf{P}_{\mathbf{A}}\mathbf{Y}_{\perp}\rVert^{2}_{F}\right]\geq 0.}$ (51)
For comparison, we also consider the performance of a two-layer neural network model trained with random weights in the first layer (i.e., Random Features), such that $\mathbf{W}_{\mathrm{random}}\in\mathbb{R}^{m_{l}\times m_{l}}$ has elements sampled from $\mathcal{N}(0,1)$. The hidden activities are given by
$\mathbf{A}_{\mathrm{random}}=\operatorname{ReLU}\left(\mathbf{X}\mathbf{W}_{\mathrm{random}}\right).$ (52)
Since $\mathbf{W}_{\mathrm{random}}$ is independent of $\mathbf{Y}$, the above hidden representation lacks alignment with the task-relevant structure that Forward Projection encodes due to the $\mathbf{YU}$ term. Hence, we cannot assume $\operatorname{col}(\hat{\mathbf{Y}}_{\mathrm{linear}})\subset\operatorname{col}(\mathbf{A}_{\mathrm{random}})$. Lastly, due to the random selection of $\mathbf{W}_{\mathrm{random}}$, we have
$\mathbb{P}\left[(\mathbf{A}_{\mathrm{random}})_{ij}=0\right]=\mathbb{P}\left[(\mathbf{X}\mathbf{W}_{\mathrm{random}})_{ij}\leq 0\right]=\frac{1}{2}$ (53)
Therefore, $\mathbf{A}_{\mathrm{random}}$ represents a random, nonlinear transformation of $\mathbf{X}$, with equal dimension, and reduced representational capacity due to element-wise sparsity. Only the output layer $\textbf{W}_{2}$ may be trained, such that:
$\hat{\mathbf{W}}_{2}:=(\mathbf{A}_{\mathrm{random}}^{\top}\mathbf{A}_{\mathrm{random}})^{-1}\mathbf{A}_{\mathrm{random}}^{\top}\mathbf{Y}.$ (54)
Unlike $\mathbf{A}$, which is correlated with the label, there is no reason to expect, over the distribution of $\mathbf{W}_{\mathrm{random}}$, that $\mathbf{A}_{\mathrm{random}}$ is a more informative design matrix than $\mathbf{X}$. Therefore, the error of this Random Features network is not expected to improve upon $E_{\mathrm{linear}}$. We note that in the special case examined here, the hidden dimension is equal to the input dimension, a framework which contrasts with the typical usage of Random Features [23], where high-dimensional random projection is employed to generate a large, representative feature set. Unlike Forward Projection, which provides the improvement in (51), Random Features hidden layer activations do not provide such an expectation in a hidden layer of equal dimension to the input. Hence, this comparison reveals that ReLU nonlinearity alone is insufficient for expected improvement – the label-informed initialization is also essential.
This analysis compares a regression model with $m_{l}\cdot m_{L}$ parameters in a single layer trained by ordinary least squares regression to a two-layer neural network with $m_{l}(m_{l}+m_{L})$ parameters, trained by Forward Projection; and a two-layer neural network with $m_{l}$ random features and $m_{l}m_{L}$ trainable parameters. Although the Forward Projection model has the greatest number of free parameters, it is noted that these are not jointly optimised, but instead may be separated into $m_{l}$ linear regressions in $\mathbf{W}_{1}$ and, subsequently, $m_{L}$ linear regressions in $\mathbf{W}_{2}$.
Differences in the modelling of Forward Projection, Random Features and linear regression each present distinct profiles of advantages and limitations. Despite these architectural differences, this analysis establishes that Forward Projection offers a principled, backpropagation-free training method with expected improvement over linear regression – an expectation absent from Random Features of equal dimension.
@@ A.10 Iterative learning with Forward Projection
where $\mathbf{z}_{j,l}=\mathbf{a}_{j,l-1}\mathbf{W}_{l}$. The gradient of $\mathcal{L}_{l}^{(t)}$ with respect to $\mathbf{W}_{l}$ is [45]:
@@
where $\eta$ is the learning rate. Optionally, regularisation may be included via the ridge penalty [45]:
$\nabla_{\mathbf{W}_{l}}\mathcal{L}_{l}^{(t)}=\frac{2}{B}\left(\mathbf{A}_{l-1}^{(t)}\right)^{\top}\left(\mathbf{A}_{l-1}^{(t)}\mathbf{W}_{l}-\tilde{\mathbf{Z}}_{l}^{(t)}\right)+\frac{2}{B}\lambda\mathbf{W}_{l}.$ (59)
@@
A.11 Modelling high-dimensional labels
We now analyse the dependence of the expected local error on the layer and label dimensions. From (29), we have,
$\hat{\mathbf{Y}}_{l}:=g_{l}^{-1}\left(\mathbf{Z}_{l}-g_{l}(\mathbf{A}_{l-1}\mathbf{Q}_{l})\right)\mathbf{U}_{l}^{+}.$ (60)
@@ A.11 Modelling high-dimensional labels
$\hat{\mathbf{Y}}_{l}:=\left(\mathbf{Z}_{l}-\mathbf{A}_{l-1}\mathbf{Q}_{l}\right)\mathbf{U}_{l}^{+}.$ (61)
@@
Recalling that the realised potentials are estimates of pre-defined target potentials, such that $\mathbf{Z}_{l}=\tilde{\mathbf{Z}}_{l}-\mathbf{E}_{l}$, we have,
$\hat{\mathbf{Y}}_{l}:=\left(\tilde{\mathbf{Z}}_{l}-\mathbf{E}_{l}-\mathbf{A}_{l-1}\mathbf{Q}_{l}\right)\mathbf{U}_{l}^{+}.$ (62)
Applying the identity function as $g_{l}$ in equation (4), we have,
$\tilde{\mathbf{Z}}_{l}=\mathbf{A}_{l-1}\mathbf{Q}_{l}+\mathbf{Y}\mathbf{U}_{l}.$ (63)
Substituting, we have,
$\hat{\mathbf{Y}}_{l}=\left(\mathbf{Y}\mathbf{U}_{l}-\mathbf{E}_{l}\right)\mathbf{U}_{l}^{+}.$ (64)
Assuming that $\mathbf{U}$ has full row rank, and that $m_{l}\geq m_{L}$, we have $\mathbf{U}_{l}\mathbf{U}^{+}_{l}=\mathbf{I}_{m_{L}}$ [50]. Therefore,
$\hat{\mathbf{Y}}_{l}=\mathbf{Y}-\mathbf{E}_{l}\mathbf{U}_{l}^{+}.$ (65)
As the realised potential is given by $\mathbf{Z}_{l}=\mathbf{P}_{\mathbf{A}_{l-1}}\tilde{\mathbf{Z}}_{l}$, the error is given by
$\mathbf{E}_{l}=\tilde{\mathbf{Z}}_{l}-\mathbf{Z}_{l}=\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\tilde{\mathbf{Z}}_{l}=\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\left(\mathbf{A}_{l-1}\mathbf{Q}_{l}+\mathbf{Y}\mathbf{U}_{l}\right),$ (66)
where $\mathbf{P}_{\mathbf{A}_{l-1}}=\mathbf{A}_{l-1}\left(\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1}\right)^{-1}\mathbf{A}_{l-1}^{\top}$ is the projection matrix of the previous layer activity $\mathbf{A}_{l-1}$. As $\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\mathbf{A}_{l-1}=0$, we have
$\mathbf{E}_{l}=\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\mathbf{Y}\mathbf{U}_{l}.$ (67)
The squared error is given by the squared Frobenius norm
$\lVert\mathbf{E}_{l}\rVert^{2}_{F}=\operatorname{Tr}\left(\mathbf{U}_{l}^{\top}\mathbf{Y}^{\top}\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\mathbf{Y}\mathbf{U}_{l}\right).$ (68)
As $\mathbf{U}_{l}\in\mathbb{R}^{m_{L}\times m_{l}}$ has entries sampled from $\mathcal{N}(0,1)$, we have the expectation
$\mathbb{E}_{\mathbf{U}_{l}}\left[\lVert\mathbf{E}_{l}\rVert^{2}_{F}\right]=m_{l}\cdot\lVert\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\mathbf{Y}\rVert^{2}_{F}.$ (69)
Firstly, it is observed that $\mathbb{E}\left[\lVert\mathbf{E}_{l}\rVert^{2}_{F}\right]$ grows linearly with the number of columns in $\mathbf{Y}$, i.e. $m_{L}$. Secondly, it is observed that $\mathbb{E}\left[\lVert\mathbf{E}_{l}\rVert^{2}_{F}\right]$ grows linearly with the number of columns in $m_{l}$, which corresponds to equal expected error in each column of $\textbf{Z}_{l}$. Assuming $\mathbf{U}_{l}\mathbf{U}_{l}^{+}=\mathbf{I}_{m_{L}}$, we have
$\mathbf{E}_{l}\mathbf{U}_{l}^{+}=\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\mathbf{Y}\mathbf{U}_{l}\mathbf{U}_{l}^{+}=\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\mathbf{Y}$ (70)
Substituting into (65), and taking the squared frobenius norm we have the squared error of the label prediction
$\mathbb{E}_{\mathbf{Q}_{l},\mathbf{U}_{l}}\left[\lVert\mathbf{Y}-\hat{\mathbf{Y}}_{l}\rVert^{2}_{F}\right]=\mathbb{E}_{\mathbf{Q}_{l},\mathbf{U}_{l}}\left[\lVert\mathbf{E}_{l}\mathbf{U}_{l}^{+}\rVert^{2}_{F}\right]=\lVert\left(\mathbf{I}-\mathbf{P}_{\mathbf{A}_{l-1}}\right)\mathbf{Y}\rVert^{2}_{F}$ (71)
Therefore $\mathbb{E}_{\mathbf{Q}_{l},\mathbf{U}_{l}}\left[\lVert\mathbf{Y}-\hat{\mathbf{Y}}_{l}\rVert^{2}_{F}\right]$ increases with the component of $\mathbf{Y}$ orthogonal to $\operatorname{col}(\mathbf{A}_{l-1})$. As $\operatorname{rank}(\mathbf{A}_{l-1})\leq m_{l-1}$, enlarging $m_{l-1}$ may increase the capacity for $\operatorname{col}(\mathbf{A}_{l-1})$ to represent $\mathbf{Y}$.
@@ A.12 Training complexity of other methods
In this section, we describe the training complexity involved in each step of the different methods that are used in this paper for benchmarking purposes. Hereafter, $N$ refers to the sample size, $N_{e}$ to the number of training epochs, $m$ the hidden layer dimension (it is assumed all layers have dimension $m$), $m_{L}$ to label dimension and $B$ to batch size.
@@ A.12.2 Local Supervision
In local supervision, after the forward pass through the main layer ($N_{e}Nm^{2}$ MACs), an auxiliary forward pass is conducted through the local supervision head ($N_{e}Nmm_{L}$ MACs). Subsequently, the auxiliary gradient is computed at the local supervision head ($N_{e}Nmm_{L}$ MACs), which is backpropagated through the main layer ($N_{e}Nm^{2}$ MACs). Forward weight updates for the main layer occur at the end of each batch ($N_{e}B^{-1}Nm^{2}$ MACs). Optionally, the auxiliary layer may be trained, requiring gradient computation ($N_{e}Nmm_{L}$ MACs) and weight updates ($N_{e}B^{-1}Nmm_{L}$ MACs). Local supervision requires storage of main layer weights ($m^{2}$), auxiliary layer weights ($mm_{L}$), main layer input activations ($Bm$), main layer output activations ($Bm$), auxiliary layer output activations ($Bm_{L}$) and weight updates ($m^{2}$).
@@ A.12.3 Forward-Forward (FF)
For a given layer, FF requires two forward passes – for “positive” and “negative” data, respectively ($2N_{e}Nm^{2}$ MACs) [16]. A goodness metric is computed for both positive and negative activations, requiring $2N_{e}Nm$ MACs. Local gradients are computed for both positive and negative passes ($2N_{e}Nm^{2}$ MACs). Forward weight updates occur with each batch ($N_{e}B^{-1}Nm^{2}$ MACs). At inference time, forward passes are required for each possible class label ($Nm^{2}m_{L}$ MACs). Forward-Forward requires storage of weight parameters ($m^{2}$), positive and negative input activations ($2Bm$), positive and negative output activations ($2Bm$) and weight updates ($m^{2}$).
@@ A.12.4 Difference Target Propagation (DTP)
DTP replaces gradient transport with targets formed using a learned approximate inverse (decoder) and a difference correction. Training requires an encoder forward pass ($N_{e}Nm^{2}$ MACs), and two decoder passes ($2N_{e}Nm^{2}$ MACs) to perform the difference correction operation between targets and observed activations [1]. Decoder training employs a denoising style loss, forwarding through both encoder ($N_{e}Nm^{2}$ MACs) and decoder ($N_{e}Nm^{2}$ MACs), before computing the decoder weight gradient ($N_{e}Nm^{2}$ MACs). Finally, the encoder weight gradient is computed with respect to the local target loss ($N_{e}Nm^{2}$ MACs). Weight updates for both encoder and decoder occur after each batch ($2N_{e}B^{-1}Nm^{2}$ MACs). At inference time, only the encoder forward pass is required. Assuming batch size $B=1$, training requires a forward pass ($N_{e}Nm^{2}$ MACs) and $8N_{e}Nm^{2}$ further MACs per layer. DTP requires storage of encoder and decoder weight parameters ($2m^{2}$), input activations ($Bm$), output activations ($Bm$), targets ($Bm$) and weight updates ($m^{2}$), which can be performed sequentially for encoder and decoder.
@@ A.12.5 Predictive Coding
PC performs $k$ iterations of activity inference that alternate forward prediction ($kN_{e}Nm^{2}$ MACs) and feedback error projection through a decoder network ($kN_{e}Nm^{2}$ MACs), followed by a local Hebbian weight gradient computation ($N_{e}Nm^{2}$ MACs) [5]. Forward weight updates occur at the end of each batch ($N_{e}B^{-1}Nm^{2}$ MACs). Letting $k=1$, $B=1$, training requires a forward pass ($N_{e}Nm^{2}$ MACs) and $3N_{e}Nm^{2}$ further MACs per layer. Predictive coding requires storage of weight parameters ($m^{2}$), input activations ($Bm$), output activations ($Bm$), output targets ($Bm$) and hebbian weight updates ($m^{2}$).